Solve for x and y: 41x+53y=135
53x+41y=147

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

41x+53y=135
53x+41y=147
-----
Using a matrix methed I get:
x = 2
y = 1
============
Cheers,
Stan H.

Answer by mananth(16949)

41x+53y=135
53x+41y=147
..
add the two equations
94x+94y=282
divide by 94
x+y=3....................1
....
subtract the two equations
-12x+12y=-12
divide by -12
x-y = 1................2
add 1 &2
2x=4
x= 2
y=1